OpenStudy (anonymous):

A clutch consists of two disks with rotational inertias I1 = 4 kg∙m2 and I2 = 2 kg∙m2 which are mounted to rotate about a common axis. Disk 1 is initially rotating at 60 rpm and disk 2 is initially at rest. The two disks are then pressed together so that, after some slippage, they both rotate at the same angular speed. (b) How much kinetic energy, if any, was lost when the disks were engaged?

OpenStudy (anonymous):

again use momentum conservation the word here being ANGULAR MOMENTUM CONSERVATION I*60+I*0=(I+I)OMEGA fing omega find the difference in kinetic energy using K.E=1/2*I*(omega) square